Practice Question

The widths of platinum samples manufactured at a factory have a mean of 1 cm and a standard deviation of 0.4 cm. Find the z-scores that correspond to each of the following widths. Round your answers to three decimal places, if necessary.

1) 1.5 cm

z= _____

2) 0.7 cm

z=______
Answer #1

Solution :

Given that mean μ = 1 and a standard deviation σ = 0.4

1) X = 1.5

=> Z = (X - μ)/σ

= (1.5 - 1)/0.4

= 1.25

=> If X = 1.5 then the Z-score is 1.250

2) X = 0.7

=> Z = (X - μ)/σ

= (0.7 - 1)/0.4

= -0.75

=> If X = 0.7 then the Z-score is -0.750